# WebAnchor: Anchoring Agent Planning to Stabilize Long-Horizon Web Reasoning

### v1

@@ Abstract
Large Language Model(LLM)-based agents have shown strong capabilities in web information seeking, with reinforcement learning (RL) becoming a key optimization paradigm. However, planning remains a bottleneck, as existing methods struggle with long-horizon strategies. Our analysis reveals a critical phenomenon, plan anchor, where the first reasoning step disproportionately impacts downstream behavior in long-horizon web reasoning tasks. Current RL algorithms, fail to account for this by uniformly distributing rewards across the trajectory. To address this, we propose Anchor-GRPO, a two-stage RL framework that decouples planning and execution. In Stage 1, the agent optimizes its first-step planning using fine-grained rubrics derived from self-play experiences and human calibration. In Stage 2, execution is aligned with the initial plan through sparse rewards, ensuring stable and efficient tool usage. We evaluate Anchor-GRPO on four benchmarks: BrowseComp, BrowseComp-Zh, GAIA, and XBench-DeepSearch. Across models from 3B to 30B, Anchor-GRPO outperforms baseline GRPO and First-step GRPO, improving task success and tool efficiency. Notably, WebAnchor-30B achieves 46.0% pass@1 on BrowseComp and 76.4% on GAIA. Anchor-GRPO also demonstrates strong scalability, getting higher accuracy as model size and context length increase.
@@ 1 Introduction
Reinforcement Learning (RL) have significantly enhanced the capabilities of autonomous, tool‑augmented agents built on large language models (LLMs) for web information seeking (jin2025search; li2025websailor; zhang2025landscapeagenticreinforcementlearning). These agents often referred to as deep research (dr; groksearch; team2025tongyi), go beyond static retrieval and instead iteratively formulate queries, invoke external tools, and collect evidence from diverse online sources to answer complex questions.
@@
Despite recent advancements, deep research agents still struggle with long-horizon planning and maintaining strategy coherence (erdogan2025plan; qiao2025webresearcher). Increasing tool usage or managing context alone does not ensure consistent multi-step reasoning; instead, accumulated errors often lead to performance degradation (liu2025webexplorer). While existing reinforcement learning (RL) methods—such as context management, reward shaping, and entropy-based optimization—mitigate certain aspects, there are still gaps to address planning instability and the lack of a comprehensive long-term strategy (chung2025evaluating; wu2025resum; zhao2025repurposing; dong2025agenticreinforcedpolicyoptimization). A crucial insight is that the ability to maintain coherent long-horizon behavior hinges on early decisions (agentfounder2025). Recent work highlights that not all reasoning steps are equally important (bogdan2025thought). In particular, the first planning step often acts as a structural anchor—shaping exploration, tool usage, and evidence integration (dong2025emergent). Initial missteps can trigger cascading failures and destabilize the entire trajectory (sui2025meta; sinha2025illusion). Motivated by this, we identify a critical phenomenon in long-horizon web agents, which we term the plan anchor, as shown in Fig. 1(a). Our motivation experiments (Fig. 1(b)) demonstrate substantial performance drops in Avg Pass@1 across BC-ZH, BC-EN, and GAIA benchmarks due to incorrect first steps—by 28.7%, 30.9%, and 23.6%, respectively. These results underscore the importance of designing reward schemes that explicitly prioritize the impact of the first step.
@@ 2 Preliminary
We follow the standard ReAct-style agentic workflow (yao2023react), where an LLM-based agent interleaves Thought, Action, and Observation. At step $t$, the agent reads the trajectory history $\mathcal{H}_{t-1}$ and produces a reasoning trace $\tau_{t}$ and an executable action $a_{t}$, after which the environment returns an observation $o_{t}$. A $T$-step rollout is defined as:
@@
Following standard web agent designs (li2025websailor; gao2025turnsunlockinglonghorizonagentic), we define the action space with two tools for web exploration:
@@ 1. Performance of Anchor-GRPO:
Does WebAnchor outperform strong external agents, including proprietary models (e.g., OpenAI, DeepSeek-V3.1) and open-source models (e.g., R1-Searcher (song2025r1), WebSailor), as well as internal baselines such as GRPO and First-step GRPO, across challenging agentic benchmarks?
@@ 4.1.1 Benchmarks and Metrics
We evaluate our method on four challenging benchmarks for web-based information-seeking tasks: BrowseComp_en (bc_en), BrowseComp_zh (bc_zh), XBench-DeepSearch (xbench), and GAIA (mialon2023gaia).
@@ 5 Related Work
Recent Deep Research (DR) agents tackle long-horizon web reasoning through multi-step planning, iterative refinement, and web-scale evidence synthesis. Moving beyond standard RAG, works like WebThinker (Li2025webthinker) and WebResearcher (qiao2025webresearcher) integrate Large Reasoning Models (LRMs) for deep thinking and self-correction. To manage unstructured evidence, structural frameworks such as WebWeaver (li2025webweaverstructuringwebscaleevidence) employ adaptive hierarchies to preserve coherence. These advances are benchmarked by DeepResearch Bench (du2025deepresearch) and extended to multimodal settings (geng2025webwatcher)), forming a robust foundation for autonomous scientific research.
@@
Agentic RL has emerged as a key paradigm, transforming LLMs from passive sequence generators to autonomous agents capable of environmental interaction and multi-step decision-making. Foundational surveys have formalized this transition, highlighting how RL optimizes both internal reasoning and external actions of agents (zhang2025landscapeagenticreinforcementlearning). Multi-turn training frameworks like Agent-R1 (cheng2025agentr1trainingpowerfulllm) and AgentGym-RL (xi2025agentgymrltrainingllmagents) enhance long-horizon performance and tool-use capabilities. Researchers are also addressing challenges like sparse feedback and robust tool integration through novel reward structures, as seen in VerlTool (jiang2025verltoolholisticagenticreinforcement). These advancements, supported by meta-thinking frameworks such as ReMA (wan2025remalearningmetathinkllms) and verifiable reasoning models like RLVMR (zhang2025rlvmrreinforcementlearningverifiable), create a comprehensive ecosystem where RL powers tool-augmented AI systems.

### v2

@@ Abstract
Large Language Model(LLM)-based agents have shown strong capabilities in web information seeking, with reinforcement learning (RL) becoming a key optimization paradigm. However, planning remains a bottleneck, as existing methods struggle with long-horizon strategies. Our analysis reveals a critical phenomenon—plan anchor—where the first reasoning step disproportionately impacts downstream behavior in long-horizon web reasoning tasks. Current RL algorithms, fail to account for this by uniformly distributing rewards across the trajectory. To address this, we propose Anchor-GRPO, a two-stage RL framework that decouples planning and execution. In Stage 1, the agent optimizes its first-step planning using fine-grained rubrics derived from self-play experiences and human calibration. In Stage 2, execution is aligned with the initial plan through sparse rewards, ensuring stable and efficient tool usage. We evaluate Anchor-GRPO on four benchmarks: BrowseComp, BrowseComp-Zh, GAIA, and XBench-DeepSearch. Across models from 3B to 30B, Anchor-GRPO outperforms baseline GRPO and First-step GRPO, improving task success and tool efficiency. Notably, WebAnchor-30B achieves 46.0% pass@1 on BrowseComp and 76.4% on GAIA. Anchor-GRPO also demonstrates strong scalability, getting higher accuracy as model size and context length increase.
@@ 1 Introduction
Reinforcement Learning (RL) have significantly enhanced the capabilities of autonomous, tool‑augmented agents built on large language models (LLMs) for web information seeking (Jin et al., 2025; Li et al., 2025a; Zhang et al., 2025a). These agents often referred to as deep research (OpenAI, 2025; Grok Team, 2025; Team et al., 2025), go beyond static retrieval and instead iteratively formulate queries, invoke external tools, and collect evidence from diverse online sources to answer complex questions.
@@
Despite recent advancements, deep research agents still struggle with long-horizon planning and maintaining strategy coherence (Erdogan et al., 2025; Qiao et al., 2025). Increasing tool usage or managing context alone does not ensure consistent multi-step reasoning; instead, accumulated errors often lead to performance degradation (Liu et al., 2025). While existing reinforcement learning (RL) methods—such as context management, reward shaping, and entropy-based optimization—mitigate certain aspects, there are still gaps to address planning instability and the lack of a comprehensive long-term strategy (Chung et al., 2025; Wu et al., 2025; Zhao et al., 2025; Dong et al., 2025a). A crucial insight is that the ability to maintain coherent long-horizon behavior hinges on early decisions (Su et al., 2025). Recent work highlights that not all reasoning steps are equally important (Bogdan et al., 2025). In particular, the first planning step often acts as a structural anchor—shaping exploration, tool usage, and evidence integration (Dong et al., 2025b). Initial missteps can trigger cascading failures and destabilize the entire trajectory (Sui et al., 2025; Sinha et al., 2025). Motivated by this, we identify a critical phenomenon in long-horizon web agents, which we term the plan anchor, as shown in Fig. 1(a). Our motivation experiments (Fig. 1(b)) demonstrate substantial performance drops in Avg Pass@1 across BC-ZH, BC-EN, and GAIA benchmarks due to incorrect first steps—by 28.7%, 30.9%, and 23.6%, respectively. These results underscore the importance of designing reward schemes that explicitly prioritize the impact of the first step.
@@ 2 Preliminary
We follow the standard ReAct-style agentic workflow (Yao et al., 2023), where an LLM-based agent interleaves Thought, Action, and Observation. At step $t$, the agent reads the trajectory history $\mathcal{H}_{t-1}$ and produces a reasoning trace $\tau_{t}$ and an executable action $a_{t}$, after which the environment returns an observation $o_{t}$. A $T$-step rollout is defined as:
@@
Following standard web agent designs (Li et al., 2025a; Gao et al., 2025), we define the action space with two tools for web exploration:
@@ 1. Performance of Anchor-GRPO:
Does WebAnchor outperform strong external agents, including proprietary models (e.g., OpenAI, DeepSeek-V3.1) and open-source models (e.g., R1-Searcher (Song et al., 2025), WebSailor), as well as internal baselines such as GRPO and First-step GRPO, across challenging agentic benchmarks?
@@ 4.1.1 Benchmarks and Metrics
We evaluate our method on four challenging benchmarks for web-based information-seeking tasks: BrowseComp_en (Wei et al., 2025), BrowseComp_zh (Zhou et al., 2025), XBench-DeepSearch (Xbench Team, 2025), and GAIA (Mialon et al., 2023).
@@ 5 Related Work
Recent Deep Research (DR) agents tackle long-horizon web reasoning through multi-step planning, iterative refinement, and web-scale evidence synthesis. Moving beyond standard RAG, works like WebThinker (Li et al., 2025b) and WebResearcher (Qiao et al., 2025) integrate Large Reasoning Models (LRMs) for deep thinking and self-correction. To manage unstructured evidence, structural frameworks such as WebWeaver (Li et al., 2025c) employ adaptive hierarchies to preserve coherence. These advances are benchmarked by DeepResearch Bench (Du et al., 2025) and extended to multimodal settings (Geng et al., 2025)), forming a robust foundation for autonomous scientific research.
@@
Agentic RL has emerged as a key paradigm, transforming LLMs from passive sequence generators to autonomous agents capable of environmental interaction and multi-step decision-making. Foundational surveys have formalized this transition, highlighting how RL optimizes both internal reasoning and external actions of agents (Zhang et al., 2025a). Multi-turn training frameworks like Agent-R1 (Cheng et al., 2025) and AgentGym-RL (Xi et al., 2025) enhance long-horizon performance and tool-use capabilities. Researchers are also addressing challenges like sparse feedback and robust tool integration through novel reward structures, as seen in VerlTool (Jiang et al., 2025). These advancements, supported by meta-thinking frameworks such as ReMA (Wan et al., 2025) and verifiable reasoning models like RLVMR (Zhang et al., 2025b), create a comprehensive ecosystem where RL powers tool-augmented AI systems.